OpenStudy (matt6288):

Which of the following shows a factor of 100b^2 - 144a^4? A.10b - 12a B.10b+ 12a C.10b + 12a^2 D.50b - 72a^2

OpenStudy (anonymous):

what Luis said

OpenStudy (matt6288):

Thx

OpenStudy (anonymous):

C because... \[\LARGE 100b^2\longrightarrow (10b)^2 , 144a^4\longrightarrow (12a^2)^2\] So we have: \[\LARGE (10b)^2-(12a^2)^2=\left[(10b)-(12a^2)\right] \cdot \left[(10b)+(12a^2)\right]\] So the answer is C